UT May 09 2022 - DEC Night



Night Info

  • Starting Config - Bino LUCIs - LUCI-1 authorize
  • Sunset 19:06
  • Nautical Twilight Ends 20:05
  • Astro Twilight Ends 20:39
  • Moon sets 01:54 - illumination 54%
  • Astro Twilight Starts 03:52
  • Nautical Twilight Starts 04:26
  • Sunrise 05:25

Staff

  • OSA: J.Williams
  • ISp: E.Solheid
  • Mtn Manager: J.Prothro
  • ISA: B.Rothberg, R.T.Gatto
  • AO: J.C.Guerra
  • Special Guests:

Night Plan

  • LUCI-1 AO/ESM Checkout
    • General Notes
      • use SOUL latest for this test, then revert on Monday morning if any problems
      • AO_Testing_V87
      • ESM_Testing_V42
    • OSCO - 0.5hr
    • Bayside Stages Fix - IT 8602 - 0.5hr
      • use a standard on-axis AO Ref star from AO OT
      • verify bayside stages don't attempt to move and then timeout when moving out of patrol field
      • use old software as control sample for same star
      • doesn't require best conditions - if can close loop, then can run test
    • Bin 1 test - 0.5hr
      • run through a science-like script from OT on any Bin 1 star
      • take TNs
    • ESM observations of faint Bin 4 targets - 0.5hr
      • Already Generated Mixed Mode Scripts: /home/lbto/Service_Observing/2022A_D_E/LUCI/AO_ESM/Bin4_ESM_Faint_Test
      • GD 321 - R=16.6 WFSMag=17.7
        • GD321_Bin4_R16pt6_ESM_SL.xml - mixed mode; GD321_Bin4_R16pt6_L1.xml - mono
        • visible all night
        • 9pt dither H-band - 1min shutter time per dither
      • SBSS 1510+510 R=17.1 WFSMag=17.96
        • GD321_Bin4_R16pt6_ESM_SL.xml - mixed mode; SBSS1510+510_Bin4_R17pt1_L1.xml - mono
        • visible all night
        • 9pt dither H-band - 1min shutter time per dither
      • 6C 165641+474241 QSO Rmag=17 (18.2)
        • 6C165641_QSO_Bin4_R17_ESM_SL.xml - mixed mode; 6C165641_QSO_Bin4_R17_L1.xml - mono
        • visible 21:00-twilight
        • 9pt dither H-band - 1min shutter time per dither
    • Angular Separation tests - 0.5-1hr
      • in OT: AO STARS - FULL DITHERS *NEW*-->Binaries
      • May only do some
        • Target 2 BD+00 2888 R=10.7 - separation 0.225" - need 256x2048 subframe - visible twilight-01hr
          • Target 5 GD 135 R=14.6 - separation 0.3" - no subframe - visible twilight-00hr
        • Target 4 SBSS 1148+564 R=14.9 separation 0.6" - no subframe - visible twilight - 02:45hr
        • Target 6 2MASSJ1128+3136 R=13.5 - 1" separation - no subframe - visible twilight - 01:30hr
        • Target 7 2MASSJ13061537+2043444 R=8.5 - 1.6" separation -64x2048 subframe - visible twilight-twilight
          • Note: given separation and lack of GS only 1 PA script is available - companion should be visible directly below main star
          • Target 9 2MASSJ10590395+6349283 Bin 2 R=12.7 - 1.57" separation - No subframe backup - visible twilight-02:45hr
        • Target 3 BD+33 2137 R=10.3 - 2.4" separation no subframe
          • Target 10 2MASSJ15553957+4025135 Bin2/3 R=13.3 - 2.3" separation - visible 01:45-twlight
    • ESM Characterization (if time) - 1hr
      • M57 Ring Nebula - 23:30-twilight
      • M92 - 21hr-twilight
  • Reconfigure for MODS-Bino D-time Science
    • LOFAR target - Becker - 2hrs - visible until twilight
      • thin cirrus OK, seeing < 1.2", no airmass restrictions

Poor Weather Backup

  • LUCI-Bino
    • Rothberg Local Mergers
      • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V47.xml
        • 1 hour per target - 1hr15 with telluric
        • Not in order of priority - observe best placed target
        • NGC 4004 - DEC +27 - twilght-02hr
        • UGC 9829 - DEC -01 - 21:30-twilight
        • NGC 3597 - DEC -23 - twilight-22hr
    • Photometric Standards - any seeing, but must be clear
      • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards
      • Select stars based on visibility, need to get stars at two different airmasses
  • LBCs
    • Photometric Standards - any seeing, but must be clear
      • OT File: /home/lbto/Service_Observing/2022A_D_E/LBC/New_PhotFields/New_Phot_Fields_Master_OT_V1.1.xml
  • MODS
    • Spetrophotometric Standards
      • Observe all modes/configurations - any seeing, but must be clear


Night Log

Closed Dome

  • UT 00:07 - bringing up LUCIs, init, test darks
    • all OK
    • L1 0001-0002, L2 0001-0002 - 2.51 sec dark
  • UT 00:14 - putting 0.5" LS away and N30 into L1 fpu - will check fieldstop
  • UT 00:20 - waking up MODS
  • UT 02:21 - taking simSnaps
  • UT 00:22 - N30 Camera LUCI1 error - IT 8183
    • camera at 471
    • init - moved to -299 - no limit switch hit - HIRMAO shows switch 11 off switch 12 on
    • selected N30, switches 11 and 12 on in HIRAMO position is 330 on MCU
  • UT 00:30 - now checking fieldstop
    • L1 0003-0006
  • UT 00:33 - simSnaps OK - just the usually doughnuts (mmm...doughnuts....)
  • UT 01:01 - winds over limits
  • UT 02:51 - winds still over the limits - sometimes we trend as low as 28m/s
  • UT 03:20 - winds still over limits - last particle count for 25microns over the limit - fun fun fun
  • UT 04:41 - winds still over limits - steady gusts at 30m/s
  • UT 06:31 - gusts are WAY down...to 27.5 m/s
  • UT 08:15 - might be trending downwards - winds are gusting to 23m/s

OSCO

  • UT 08:55 - getting ready to open - will start with OSCO to assess
  • UT 09:03 - pointing check
  • UT 09:07 - sending collimation preset
    • star might be a double - or seeing is truly terrible
    • trying a new star a bit brighter and possibly not so doubly
    • FWHM 1.47" guider 1.22" WFS
  • FWHM settling around 1.7" - skipping OSCO

Bayside Stage Test

  • UT 09:16 - sending preset to SBSS1709+535 Rmag=12.8
  • WFSmag = 15.2
  • not going to close the loops with full modes, tip and tilt should be fine for test
  • UT 09:23 - exposure started
  • loading a quicker 5dit x 4 NDIT and skipping down to the bigger jumps - coultdn't load script with 5x4 - very weird, so continuing with original
    • Mon May 9 09:28:54.342 2022 left Offset alert for LUCIFER: Pause{FLAO}[99]
      Mon May 9 09:28:54.359 2022 left Offset alert for LUCIFER: AGw command status: AGW1: Requested position is Out of Range
      Mon May 9 09:28:54.376 2022 left Offset alert for LUCIFER: isPosAccessible: probe cannot move to requested position (SFP) (x,y): 201.068, 110.646 mm.
      Mon May 9 09:28:54.392 2022 left Offset alert for LUCIFER: probe failed moving to new position: AGW1: Requested position is Out of Range
      Mon May 9 09:28:54.409 2022 left Offset alert for LUCIFER: OffsetXY {FLAO}[111] - [AOException] RETRY: Offset outside FoV limits. Target xyz is -137.28 7.67 52.89 (-5002) VALUE_OUT_OF_RANGE_ERROR
      Mon May 9 09:28:54.426 2022 left Offset command complete with warnings for LUCIFER, Sound=offsetWarning.wav
  • UT 09:32 - 125",100" offset
    • Mon May 9 09:32:02.145 2022 left Offset alert for LUCIFER: Pause{FLAO}[99]
      Mon May 9 09:32:02.163 2022 left Offset alert for LUCIFER: AGw command status: AGW1: Requested position is Out of Range
      Mon May 9 09:32:02.179 2022 left Offset alert for LUCIFER: isPosAccessible: probe cannot move to requested position (SFP) (x,y): 216.112, 110.677 mm.
      Mon May 9 09:32:02.195 2022 left Offset alert for LUCIFER: probe failed moving to new position: AGW1: Requested position is Out of Range
      Mon May 9 09:32:02.211 2022 left Offset alert for LUCIFER: OffsetXY {FLAO}[116] - [AOException] RETRY: Offset outside FoV limits. Target xyz is -152.24 7.61 52.89 (-5002) VALUE_OUT_OF_RANGE_ERROR
      Mon May 9 09:32:02.228 2022 left Offset command complete with warnings for LUCIFER, Sound=offsetWarning.wav
  • UT 09:34 100",125" offset
    • Mon May 9 09:35:11.437 2022 left Offset alert for LUCIFER: Pause{FLAO}[99]
      Mon May 9 09:35:11.448 2022 left Offset alert for LUCIFER: AGw command status: AGW1: Requested position is Out of Range
      Mon May 9 09:35:11.467 2022 left Offset alert for LUCIFER: isPosAccessible: probe cannot move to requested position (SFP) (x,y): 201.093, 125.661 mm.
      Mon May 9 09:35:11.484 2022 left Offset alert for LUCIFER: probe failed moving to new position: AGW1: Requested position is Out of Range
      Mon May 9 09:35:11.501 2022 left Offset alert for LUCIFER: OffsetXY {FLAO}[121] - [AOException] RETRY: Offset outside FoV limits. Target xyz is -137.35 22.65 52.89 (-5002) VALUE_OUT_OF_RANGE_ERROR
      Mon May 9 09:35:11.519 2022 left Offset command complete with warnings for LUCIFER, Sound=offsetWarning.wav
  • UT 09:37 - DIMM shows 1.28"-1.3" - improving?
  • UT 09:37 - offset
    • Mon May 9 09:38:22.939 2022 left Offset alert for LUCIFER: Pause{FLAO}[99]
      Mon May 9 09:38:22.954 2022 left Offset alert for LUCIFER: AGw command status: AGW1: Requested position is Out of Range
      Mon May 9 09:38:22.970 2022 left Offset alert for LUCIFER: isPosAccessible: probe cannot move to requested position (SFP) (x,y): 201.091, 125.665 mm.
      Mon May 9 09:38:22.990 2022 left Offset alert for LUCIFER: probe failed moving to new position: AGW1: Requested position is Out of Range
      Mon May 9 09:38:23.008 2022 left Offset alert for LUCIFER: OffsetXY {FLAO}[126] - [AOException] RETRY: Offset outside FoV limits. Target xyz is -137.35 22.65 52.89 (-5002) VALUE_OUT_OF_RANGE_ERROR
      Mon May 9 09:38:23.025 2022 left Offset command complete with warnings for LUCIFER, Sound=offsetWarning.wav
  • UT 09:41
    • Mon May 9 09:41:37.067 2022 left Offset alert for LUCIFER: Pause{FLAO}[99]
      Mon May 9 09:41:37.082 2022 left Offset alert for LUCIFER: AGw command status: AGW1: Requested position is Out of Range
      Mon May 9 09:41:37.099 2022 left Offset alert for LUCIFER: isPosAccessible: probe cannot move to requested position (SFP) (x,y): 216.138, 125.7 mm.
      Mon May 9 09:41:37.115 2022 left Offset alert for LUCIFER: probe failed moving to new position: AGW1: Requested position is Out of Range
      Mon May 9 09:41:37.132 2022 left Offset alert for LUCIFER: OffsetXY {FLAO}[131] - [AOException] RETRY: Offset outside FoV limits. Target xyz is -152.32 22.59 52.89 (-5002) VALUE_OUT_OF_RANGE_ERROR
      Mon May 9 09:41:37.148 2022 left Offset command complete with warnings for LUCIFER, Sound=offsetWarning.wav
  • Condtions seem steadier will try OSCO

OSCO Deux

  • UT 09:46 - sending preset
  • HD 168811 Rmag=7.54
  • WFSMag =
  • Thought it was steadier, was for 10min or so, now unsteady - will still try
  • UT 09:53 - failed to optimize - ASM failure - RIP'd
  • UT 09:58 - recovered - sending again - guider shows 1.3"
  • safeskipping, H02 problems
  • 1st frame no NCPA
    • L1 0020-0029
  • UT 10:08 - failed on 5",5" offset
    • loop opened - don't know why
    • 2022.05.09 10:08:57 s | user medium | luci.luc | LBT | TelescopeServer | CommandThread.java#run(479) | left side DETXY ABS Offset sent to IIF: dx =-5.000 dy=-5.000 arcsec, dTheta=0.000
      2022.05.09 10:08:58 s | error high | luci.luc | LBT | TelescopeServer | RMITelescopeImpl.java#iifCommandMessage(1439) | IIF OffsetPointing command for left side(s) failed! 0- OffsetPointing result status: Error 1- Pause{FLAO}[99] - Illegal command for state AOArbitrator.Ready
      2022.05.09 10:08:58 s | error high | luci.luc | LBT | TelescopeServer | RMITelescopeImpl.java#commitError(1012) | commit failed exception de.rub.astro.lucifer.control.telescope.CommitFailedException: IIF command failed
      Left command:OffsetPointing failed.
      OK
  • trying offset again - failed
    • 2022.05.09 10:10:13 s | error high | luci.luc | LBT | TelescopeServer | RMITelescopeImpl.java#iifCommandMessage(1439) | IIF OffsetPointing command for left side(s) failed! 0- OffsetPointing result status: Error 1- Pause{FLAO}[99] - Illegal command for state AOArbitrator.Ready
    • This failed because the loop was opened - miscomunication
  • NCPA command caused error?
  • getting skipframes with NCPA
  • Why did the man put a sweater on his hot dog? Because it was a chili dog. Wocka Wocka Wocka!
  • UT 10:12 - continuing with NCAP off --- and bonk it failed
    • 2022.05.09 10:12:51 s | error high | luci.luc | LBT | TelescopeServer | RMITelescopeImpl.java#iifCommandMessage(1439) | IIF OffsetPointing command for left side(s) failed! 0- OffsetPointing result status: Error 1- Pause{FLAO}[99] - Illegal command for state AOArbitrator.Ready
      2022.05.09 10:12:51 s | error high | luci.luc | LBT | TelescopeServer | RMITelescopeImpl.java#commitError(1012) | commit failed exception de.rub.astro.lucifer.control.telescope.CommitFailedException: IIF command failed
      Left command:OffsetPointing failed.
      OK
    • JC Guerra says pause failed
  • UT 10:13 - sending from offset in script again - failed again
    • 2022.05.09 10:14:04 s | error high | luci.luc | LBT | TelescopeServer | RMITelescopeImpl.java#iifCommandMessage(1439) | IIF OffsetPointing command for left side(s) failed! 0- OffsetPointing result status: Error 1- Pause{FLAO}[99] - Illegal command for state AOArbitrator.Ready
      2022.05.09 10:14:04 s | error high | luci.luc | LBT | TelescopeServer | RMITelescopeImpl.java#commitError(1012) | commit failed exception de.rub.astro.lucifer.control.telescope.CommitFailedException: IIF command failed
      Left command:OffsetPointing failed.
      OK
  • IIF says loop is closed,
    • From IIF:
    • Mon May 9 10:14:03.177 2022 left Offset command started for LUCIFER
      Mon May 9 10:14:04.208 2022 left Offset alert for LUCIFER: Pause{FLAO}[99] - Illegal command for state AOArbitrator.Ready
      Mon May 9 10:14:04.229 2022 left Offset command failed for LUCIFER, Sound=offsetFailed.wav
  • Trying to parse what happened ... loops opened because the max # of skipframes occurred before the offset was triggered (3x in a row?)
  • UT 10:21 - ASM back
  • UT 10:22 - resending preset , 1.5" on DIMM
    • L1 0029 - no NCPA
    • L1 0030-0034 NCPA on
    • L1 0035-0048 NCPA off
  • NCPA is triggering skipframes - smoking gun?
  • w/NCPA off 1-5% skipframes
  • w/NCPA on 15-75% skipframes

SBSS 1709+535

  • UT 10:36 - sending preset
  • Plan is to close the loop in Bin 2 - see if NCPA on triggers skipframes in similar conditions
  • Rmag=12.8, WFSmag = 14.8
  • Did I tell you the one about the man with the light bulb in his nose? He was lightheaded! Get it? Wocka Wocka Wocka!
  • JC Guerra says for optimize skipping half the tests - says this was what was done with OSCO to compare
  • UT 10:46 - DIMM 1.2"
  • FWHM ~-0.12"
  • UT 10:49 - bayside stage was retrying - was in timeout - but the offset was -1",0" - timeout was 40seconds
  • was working fine before on same script - same target - is there a hardware problem on bayside stages?
  • different rotator position than before - could this be an issue?
  • Seems like we are testing the bayside stage offset again - two for one!
  • UT 10:55 - skipframes
  • UT 10:56 - loops opened due to skipframes
    • L1 0049-0060 NCPA Off
  • UT 10:57 - going back to earlier in script - loops openeing
  • UT 10:58 - gain is low, testing small offsets, NCPA is off
  • One issue found - the bayside stage timeout was 40sec and then eventually was able to do the offset (within the AO patrol field)

PG1530+212

  • UT 11:05 - sending preset
  • Rmag=11.8 - should end up as Bin 2 (hopefully)
  • UT 11:13 - OCAM went not live
  • UT 11:17 - OCAM still not live
  • We are DONE

Closing

  • UT 11:24 - close up - put LUCI-1 in safe config
  • UT 11:37 - putting MODS to bed

-- BarryRothberg - 06 May 2022
Topic revision: r11 - 09 May 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback